# Proximal Interacting Particle Langevin Algorithms

### v2

@@ 3.2 Proximal Particle Gradient Descent Methods
$\mathrm{d}\mathbf{X}_{t}^{i,N}$ $=-\nabla_{x}U^{\lambda}(\bm{\theta}_{t}^{N},\mathbf{X}_{t}^{i,N})\mathrm{d}t+\sqrt{2}\mathrm{d}\mathbf{B}_{t}^{i,N},$
where $i=1,\ldots,N$ and $(\mathbf{B}_{t}^{i,N})_{t\geq 0}$ is a family of independent Brownian motions. Discretising this SDE system, we obtain similar algorithms to PIPULA and MYIPLA without the term $\sqrt{2\gamma/N}\xi_{n+1}^{0,N}$ in (11) and (13), which can be seen as proximal versions of PGD [Kuntz et al., 2023]. Accordingly, we term these methods as proximal PGD (PPGD) and Moreau-Yosida PGD (MYPGD), respectively. We provide a detailed description of these methods and their theoretical analysis in Appendix B.
@@ 4 Nonasymptotic analysis
In this section, we provide theoretical analysis of the parameter estimates obtained by the proximal interacting particle algorithms. We first introduce the assumptions that will be central for our results. Some further assumptions will be introduced later to refine some of the results.
@@ 4.1 Assumptions
It is important to note that the assumptions made for the different algorithms are similar and allow a fair comparison between the derived convergence rates. Let $g_{1},g_{2}:\mathbb{R}^{d_{\theta}}\times\mathbb{R}^{d_{x}}\to\mathbb{R}$ and recall that $U(\theta,x)=g_{1}(\theta,x)+g_{2}(\theta,x)$.
@@
The requirements on $g_{2}$ guarantee that $\operatorname{prox}_{g_{2}}^{\lambda}$ is well defined, and similarly for $g_{1}$ and its gradient. A1 results in $\nabla U^{\lambda}$ being Lipschitz in both variables with constant $L\leq L_{g_{1}}+\lambda^{-1}$ [Durmus et al., 2018, Proposition 1].
@@
There exist a constant $H>0$ such that the initial condition $Z_{0}^{N}=(\theta_{0},N^{-1/2}X_{0}^{1,N},\dots,$ $N^{-1/2}X_{0}^{N,N})$ satisfies $\mathbb{E}[\|Z_{0}^{N}\|^{2}]\leq H$.
@@
$g_{1}$ is $\mu$-strongly convex, that is, for all $v=(\theta,x),v^{\prime}=(\theta^{\prime},x^{\prime})$
$\langle v-v^{\prime},\nabla g_{1}(v)-\nabla g_{1}(v^{\prime})\rangle\geq\mu\|v-v^{\prime}\|^{2}$
for some $\mu>0$.
Remark 4.1.
Observe that $\nabla g_{2}^{\lambda}(v)=(v-\operatorname{prox}_{g_{2}}^{\lambda}(v))/\lambda$ and $\operatorname{prox}_{g_{2}}^{\lambda}$ is firmly non expansive [Durmus et al., 2018, Eq. (7)] which implies Lipschitzness $\|\operatorname{prox}_{g_{2}}^{\lambda}(v)-\operatorname{prox}_{g_{2}}^{\lambda}(v^{\prime})\|\leq\|v-v^{\prime}\|$, and we have, by the Cauchy-Schwarz inequality,
$\langle v-v^{\prime},\nabla g_{2}^{\lambda}(v)-\nabla g_{2}^{\lambda}(v^{\prime})\rangle$ $=\frac{1}{\lambda}(\|v-v^{\prime}\|^{2}-\langle v-v^{\prime},\operatorname{prox}_{g_{2}}^{\lambda}(v)-\operatorname{prox}_{g_{2}}^{\lambda}(v^{\prime})\rangle)$
@@
Therefore, under A4, $\nabla U^{\lambda}$ is also $\mu$-strongly convex
$\langle v-v^{\prime},\nabla U^{\lambda}(v)-\nabla U^{\lambda}(v^{\prime})\rangle$ $=\langle v-v^{\prime},\nabla g_{1}(v)-\nabla g_{1}(v^{\prime})\rangle+\langle v-v^{\prime},\nabla g_{2}^{\lambda}(v)-\nabla g_{2}^{\lambda}(v^{\prime})\rangle$
$\geq\mu\|v-v^{\prime}\|^{2}.$
$\lozenge$
Remark 4.2.
Let $\Omega\subset\mathbb{R}^{d_{\theta}}\times\mathbb{R}^{d_{x}}$ denote the (nonempty) set where $g_{2}$ is twice differentiable. Theorem 25.5 in Rockafellar [1970], guarantees that if $g_{2}$ is a proper, convex function, then $g_{2}$ is differentiable except in a set of measure zero, i.e., $\Omega^{\mathsf{c}}$ has measure zero. Also, by Alexandrov’s Theorem [Rockafellar, 1999], $g_{2}$ is twice differentiable almost everywhere—in particular, these points form a subset of $\textnormal{{dom}}(\nabla g_{2})$. In addition, the matrix $\nabla^{2}g_{2}$ (or alternatively its distributional counterpart, $D^{2}g_{2}$, if $\nabla^{2}g_{2}$ does not exist) is symmetric and positive definite [Alberti and Ambrosio, 1999, Proposition 7.11]. Thus, we do not need to assume that $g_{2}$ is twice differentiable everywhere since the regularity of $g_{2}$ in A1 guarantees that $g_{2}$ is differentiable except in a set of measure zero. $\lozenge$
Let $\bar{\theta}_{\star}$ be the maximiser of $p_{\theta}(y)$. Let $\mathsf{m}_{\bar{\theta}_{\star}}$ be the restriction of the Lebesgue measure $\mathsf{m}$ on $\mathbb{R}^{d_{\theta}}\times\mathbb{R}^{d_{x}}$ to the set $\{\bar{\theta}_{\star}\}\times\mathbb{R}^{d_{x}}$, which is well defined (see, e.g., Bogachev [2007, Section 10.6]).
@@
$\mathbb{E}_{X}[\|\nabla_{\theta}U(\bar{\theta}_{\star},X)\|]\leq A\quad\quad\textnormal{ and }\quad\quad\mathbb{E}_{X}[\|\nabla^{2}_{(\theta,x)}g_{2}(\bar{\theta}_{\star},X)\nabla_{(\theta,x)}g_{2}(\bar{\theta}_{\star},X)\|]\leq B$
where $X\sim\rho_{\bar{\theta}_{\star}}(x)$ with $\rho_{\bar{\theta}_{\star}}(x)\propto\exp(-U^{\lambda}(\bar{\theta}_{\star},x))$.
@@ 4.2 The proof strategy
Recall that our aim is to find the MMLE:
$\bar{\theta}_{\star}=\operatorname*{arg\,max}_{\theta}p_{\theta}(y),$
where $p_{\theta}(y)=\int e^{-U(\theta,x)}\mathrm{d}x$. Therefore, we want to provide an upper bound on the distance between the iterates of our algorithms and $\bar{\theta}_{\star}$, that is, $\mathbb{E}[\|\theta_{n}-\bar{\theta}_{\star}\|^{2}]^{1/2}$.
@@
$\pi^{N}_{\lambda,\Theta}(\theta)\propto\int_{{}^{d_{x}}}...\int_{{}^{d_{x}}}e^{-\sum^{N}_{i=1}U^{\lambda}(\theta,x_{i})}\mathrm{d}x_{1}\mathrm{d}x_{2}\dots\mathrm{d}x_{N}=\biggr{(}\int_{{}^{d_{x}}}e^{-U^{\lambda}(\theta,x)}\mathrm{d}x\biggr{)}^{N}.$
Using $\mathbb{E}[\|\theta_{n}^{N}-\bar{\theta}_{\star}\|^{2}]^{1/2}=W_{2}(\delta_{\bar{\theta}_{\star}},\mathcal{L}(\theta_{n}^{N}))$ and the fact that $W_{2}$ is a metric, the following triangle inequality holds:
$W_{2}(\delta_{\bar{\theta}_{\star}},\mathcal{L}(\theta_{n}^{N}))\leq\underbrace{W_{2}(\delta_{\bar{\theta}_{\star}},\pi_{\lambda,\Theta}^{N})}_{\textnormal{concentration}}+\underbrace{W_{2}(\pi_{\lambda,\Theta}^{N},\mathcal{L}(\bm{\theta}_{n\gamma}^{N}))}_{\textnormal{convergence}}+\underbrace{W_{2}(\mathcal{L}(\theta_{n}^{N}),\mathcal{L}(\bm{\theta}_{n\gamma}^{N}))}_{\textnormal{discretisation}}.$ (19)
The first term is a concentration result, characterising the concentration of the $\theta$-marginal of the target measure of the SDE (9)–(10) around the maximiser of $p_{\theta}(y)$. Handling this term is not trivial since the maximisers of $p_{\theta}(y)$ and of $p_{\theta}^{\lambda}(y):=\int_{{}^{d_{x}}}p_{\theta}^{\lambda}(x,y)\mathrm{d}x$ are not necessarily the same as we clarify in the next section. The second term is a convergence term, which characterises the convergence of the solution of the SDE to its target measure. The last term is a discretisation term, which characterises the error introduced by discretising the SDE. By separately bounding each term, we provide nonasymptotic results for the convergence of the proximal interacting particle algorithms. Proofs of the following results are provided in Appendices A and B.
@@ 4.3 Nonasymptotic analysis of MYIPLA
Let A1A5 hold. Let $\theta_{n}^{N}$ denote the law of the iterate (13) and $\bar{\theta}_{\star}$ be the maximiser of $p_{\theta}(y)$. Fix $\gamma_{0}\in(0,\min\{(L_{g_{1}}+\lambda^{-1})^{-1},2\mu^{-1}\})$. Then for every $\lambda>0$ and $\gamma\in(0,\gamma_{0}]$, one has
@@
Proof.
See Appendix A.2. ∎
While the full proof is given in Appendix A.2, let us unpack this result. As stated in the proof strategy, we split the errors into three terms. The first term is the concentration term in (19), which in this case can be bounded as
$W_{2}(\delta_{\bar{\theta}_{\star}},\pi_{\lambda,\Theta}^{N})\leq\|\bar{\theta}_{\star}-\bar{\theta}_{\star,\lambda}\|+W_{2}(\delta_{\bar{\theta}_{\star,\lambda}},\pi_{\lambda,\Theta}^{N}).$
The first term here can be bounded as
@@
with $A,B$ given in A5. This term quantifies the distance between maximisers of $p_{\theta}(y)$ and $p^{\lambda}_{\theta}(y)$. Next, consider the stationary measure of the SDE (9)–(10) denoted by $\pi_{\lambda,\star}^{N}(\theta,x_{1},\ldots,x_{N})$. We are interested in the $\theta$-marginal of this measure, denoted by $\pi_{\lambda,\Theta}^{N}$. Using a form of the Prékopa-Leindler inequality for strong convexity [Saumard and Wellner, 2014, Theorem 3.8], $\pi_{\lambda,\Theta}^{N}$ is $N\mu$-strongly log-concave and by Lemma A.8 of Altschuler and Chewi [2023]
$W_{2}(\pi_{\lambda,\Theta}^{N},\delta_{\bar{\theta}_{\star}})\leq\sqrt{\frac{d_{\theta}}{N\mu}},$
which concludes the bound of the concentration term. The second challenge is the convergence term in (19) which is characterised by the exponential decay of the Wasserstein distance between the $\theta$-marginal of the solution of the SDE and its stationary measure, i.e., we have
$W_{2}(\pi_{\lambda,\Theta}^{N},\mathcal{L}(\bm{\theta}_{n\gamma}^{N}))\leq e^{-\mu n\gamma}\bigg{(}\mathbb{E}[\|Z_{0}^{N}-z_{\star}\|^{2}]^{1/2}+\Big{(}\frac{d_{x}N+d_{\theta}}{N\mu}\Big{)}^{1/2}\bigg{)}.$
Finally, we have the discretisation term in (19) which is characterised by the error introduced by discretising the SDE. This term is bounded by
@@
Merely summing these terms, we obtain the final bound for the convergence of MYIPLA as given in Theorem 4.1.
@@ 4.4 Nonasymptotic analysis of PIPGLA
To derive a convergence rate for PIPGLA we make the following additional assumption.
@@
This assumption is similar to Salim et al. [2019, Assumption 5] without involving additional random variables.
@@
Let A1A5 and C1 hold. Let $\theta_{n}^{N}$ denote the law of the iterate (17) and $\bar{\theta}_{\star}$ be the maximiser of $p_{\theta}(y)$. Then for $\gamma\leq 1/L_{g_{1}}$ and $\gamma\leq\lambda\leq\gamma/(1-\mu\gamma)$, the following holds
$\mathbb{E}\left[\|\theta_{n}^{N}-\bar{\theta}_{\star}\|^{2}\right]^{1/2}\leq$ $\sqrt{\frac{d_{\theta}}{N\mu}}+\frac{\lambda^{n/2}(1-\gamma\mu)^{n/2}}{\gamma^{n/2}}W_{2}(\mathcal{L}({Z_{0}^{N}}),\pi^{N})$ (20)
$+\left(\frac{\lambda(2\gamma L_{g_{1}}(d_{\theta}+Nd_{x})+\lambda NC)}{N\left(1-\lambda(1-\mu\gamma)/\gamma\right)}\right)^{1/2},$
@@
Proof.
See Appendix A.4. ∎
Similarly to the previous result, we can split the errors as follows:
@@
Using a similar result to Altschuler and Chewi [2023, Lemma 8], the concentration term $W_{2}(\delta_{\bar{\theta}_{\star}},\pi_{\Theta}^{N})$ is bounded by
$W_{2}(\delta_{\bar{\theta}_{\star}},\pi_{\Theta}^{N})\leq\sqrt{\frac{d_{\theta}}{N\mu}}.$
We derive novel nonasymptotic bounds for the other error term (see Corollary A.6 in Appendix A.4). Intuitively, $W_{2}(\pi_{\Theta}^{N},\mathcal{L}({\theta_{n}^{N}}))$ controls the convergence to $\pi_{\Theta}^{N}$ and the error due to time discretisation.

### v3

@@ 3.2 Proximal Particle Gradient Descent Methods (PPGD)
$\mathrm{d}\mathbf{X}_{t}^{i,N}$ $=-\nabla_{x}U^{\lambda}(\bm{\theta}_{t}^{N},\mathbf{X}_{t}^{i,N})\mathrm{d}t+\sqrt{2}\mathrm{d}\mathbf{B}_{t}^{i,N}.$
Discretising this SDE system, we obtain similar algorithms to MYIPLA and PIPULA without the term $\sqrt{2\gamma/N}\xi_{n+1}^{0,N}$ in (9), which can be seen as proximal versions of PGD. Accordingly, we term these methods as proximal PGD (PPGD) and Moreau-Yosida PGD (MYPGD), respectively. We provide a detailed description of these methods and their theoretical analysis in Appendix B.
@@ 4 Nonasymptotic analysis
This section presents a theoretical analysis of the parameter estimates obtained by the PIPLA family. It is important to note that similar assumptions across algorithms enable a fair comparison of convergence rates.
@@ 4.1 Assumptions
Let $g_{1},g_{2}:\mathbb{R}^{d_{\theta}}\times\mathbb{R}^{d_{x}}\to\mathbb{R}$ and $U=g_{1}+g_{2}$.
@@
$g_{1}$ is continuously differentiable, convex, $L_{g_{1}}$-smooth and lower bounded, and $g_{2}$ is proper, convex, lower semi-continuous and lower bounded.
A1 guarantees that $\operatorname{prox}_{g_{2}}^{\lambda}$ is well defined and that $\nabla U^{\lambda}$ is Lipschitz in both variables with constant $L\leq L_{g_{1}}+\lambda^{-1}$ [Durmus et al., 2018, Proposition 1].
@@
The initial condition $Z_{0}^{N}=(\theta_{0},N^{-1/2}X_{0}^{1,N},\dots,$ $N^{-1/2}X_{0}^{N,N})$ satisfies $\mathbb{E}[\|Z_{0}^{N}\|^{2}]\leq H$ for $H>0$.
@@
$g_{1}$ is $\mu$-strongly convex.
Remark 2.
Let $v=(\theta,x)$ and $v^{\prime}=(\theta^{\prime},x^{\prime})$. We have that $\nabla g_{2}^{\lambda}(v)=(v-\operatorname{prox}_{g_{2}}^{\lambda}(v))/\lambda$ and $\operatorname{prox}_{g_{2}}^{\lambda}$ is firmly non expansive [Durmus et al., 2018, Eq. (7)] which implies Lipschitzness. By the Cauchy-Schwarz inequality,
$\langle v-v^{\prime},\nabla g_{2}^{\lambda}(v)-\nabla g_{2}^{\lambda}(v^{\prime})\rangle$
@@
Therefore, under A4, $\nabla U^{\lambda}$ is also $\mu$-strongly convex.
Remark 3.
Let $\Omega\subset\mathbb{R}^{d_{\theta}}\times\mathbb{R}^{d_{x}}$ denote the (nonempty) set where $g_{2}$ is twice differentiable. Theorem 25.5 in [Rockafellar, 1970], guarantees that if $g_{2}$ is a proper, convex function, then $g_{2}$ is differentiable except in a set of measure zero, i.e., $\Omega^{\mathsf{c}}$ has measure zero. Also, by Alexandrov’s Theorem [Rockafellar, 1999], $g_{2}$ is twice differentiable almost everywhere—in particular, these points form a subset of $\textnormal{{dom}}(\nabla g_{2})$. In addition, the Hessian $\nabla^{2}g_{2}$ (or alternatively its distributional counterpart, $D^{2}g_{2}$, if $\nabla^{2}g_{2}$ does not exist) is symmetric and positive definite [Alberti and Ambrosio, 1999, Proposition 7.11].
Let $\bar{\theta}_{\star}$ be the maximiser of $p_{\theta}(y)$. Let $\mathsf{m}_{\bar{\theta}_{\star}}$ be the restriction of the Lebesgue measure $\mathsf{m}$ on $\mathbb{R}^{d_{\theta}}\times\mathbb{R}^{d_{x}}$ to the set $\{\bar{\theta}_{\star}\}\times\mathbb{R}^{d_{x}}$, which is well defined (see, e.g., [Bogachev, 2007, Section 10.6]).
@@
$\mathbb{E}_{X}[\|\nabla_{\theta}U(\bar{\theta}_{\star},X)\|]\leq A,$
$\mathbb{E}_{X}[\|\nabla^{2}_{(\theta,x)}g_{2}(\bar{\theta}_{\star},X)\nabla_{(\theta,x)}g_{2}(\bar{\theta}_{\star},X)\|]\leq B$
where $X\sim\rho_{\bar{\theta}_{\star}}(x)$ with $\rho_{\bar{\theta}_{\star}}(x)\propto e^{-U^{\lambda}(\bar{\theta}_{\star},x)}$.
@@ 4.2 The proof strategy
Our objective is to find the MMLE $\bar{\theta}_{\star}=\operatorname*{arg\,max}_{\theta}p_{\theta}(y)$, where $p_{\theta}(y)=\int e^{-U(\theta,x)}\mathrm{d}x$. Therefore, we provide an upper bound on the distance between the iterates of our algorithms and $\bar{\theta}_{\star}$, that is, $\mathbb{E}[\|\bar{\theta}_{\star}-\theta_{n}\|^{2}]^{1/2}$.
@@
$\pi^{N}_{\lambda,\Theta}(\theta)$ $\propto\biggr{(}\int_{\mathbb{R}^{d_{x}}}e^{-U^{\lambda}(\theta,x)}\mathrm{d}x\biggr{)}^{N}.$ (15)
Note that $\mathbb{E}[\|\bar{\theta}_{\star}-\theta_{n}^{N}\|^{2}]^{1/2}=W_{2}(\delta_{\bar{\theta}_{\star}},\mathcal{L}(\theta_{n}^{N}))$. Applying the triangle inequality of the $W_{2}$ metric, it follows:
$W_{2}(\delta_{\bar{\theta}_{\star}},\mathcal{L}(\theta_{n}^{N}))\leq$ $\underbrace{W_{2}(\delta_{\bar{\theta}_{\star}},\pi_{\lambda,\Theta}^{N})}_{\textnormal{concentration}}+\underbrace{W_{2}(\pi_{\lambda,\Theta}^{N},\mathcal{L}(\bm{\theta}_{n\gamma}^{N}))}_{\textnormal{convergence}}$
$+\underbrace{W_{2}(\mathcal{L}(\theta_{n}^{N}),\mathcal{L}(\bm{\theta}_{n\gamma}^{N}))}_{\textnormal{discretisation}}.$ (16)
The concentration term characterises the concentration of the $\theta$-marginal of the target measure of the SDE (7)–(8), given by (15), around the maximiser of $p_{\theta}(y)$, denoted by $\bar{\theta}_{\star}$. Handling this term is not trivial since the maximisers of $p_{\theta}(y)$ and of $p_{\theta}^{\lambda}(y):=\int_{\mathbb{R}^{d_{x}}}p_{\theta}^{\lambda}(x,y)\mathrm{d}x$ are not necessarily the same as we clarify in the next section. The convergence term captures the convergence of the solution of the SDE to its target measure (15). Finally, the discretisation term characterises the error introduced by discretising the SDE. We provide nonasymptotic results for the convergence of the PIPLA family, proofs of the results are provided in Appendices A and B.
@@ 4.3 Nonasymptotic analysis of MYIPLA
Let A1A5 hold. Let $\theta_{n}^{N}$ denote the iterate (9) and $\bar{\theta}_{\star}$ be the maximiser of $p_{\theta}(y)$. Fix $\gamma_{0}\in(0,\min\{(L_{g_{1}}+\lambda^{-1})^{-1},2\mu^{-1}\})$. Then for every $\lambda>0$ and $\gamma\in(0,\gamma_{0}]$, one has
@@
See Appendix A.2 for the full proof. Below, we provide a sketch of the proof, following the error decomposition in (4.2). The concentration term can be split into two parts: (1) the distance between the MMLE of the original distribution, $\bar{\theta}_{\star}$, and the MMLE of the MY approximation, $\bar{\theta}_{\star,\lambda}$, and (2) the concentration of the MMLE of the MY approximation around the target regularised marginal measure $\pi_{\lambda,\Theta}^{N}$ provided in (15). This results in
$W_{2}(\delta_{\bar{\theta}_{\star}},\pi_{\lambda,\Theta}^{N})\leq\|\bar{\theta}_{\star}-\bar{\theta}_{\star,\lambda}\|+W_{2}(\delta_{\bar{\theta}_{\star,\lambda}},\pi_{\lambda,\Theta}^{N}),$
where $\bar{\theta}_{\lambda,\star}$ denotes the maximiser of $p_{\theta}^{\lambda}(y)$. We derive in Proposition A.1 a novel bound for the distance between the maximisers $\bar{\theta}_{\star},\bar{\theta}_{\star,\lambda}$, given by
@@
with $A,B$ given in A5. We observe that stronger regularisation (i.e., larger values of $\lambda$) increases the distance between the two MMLEs—intuitively, in the limit $\lambda\to\infty$, all points collapse to the minimiser of the potential $U$, therefore this leads to a larger difference between $\bar{\theta}_{\star}$ and $\bar{\theta}_{\star,\lambda}$.
Next, consider the stationary measure of the SDE (7)–(8), denoted by $\pi_{\lambda,\star}^{N}(\theta,x_{1},\ldots,x_{N})$, and its $\theta$-marginal given by $\pi_{\lambda,\Theta}^{N}$. Using a form of the Prékopa-Leindler inequality for strong convexity [Saumard and Wellner, 2014, Theorem 3.8], $\pi_{\lambda,\Theta}^{N}$ is $N\mu$-strongly log-concave and by [Altschuler and Chewi, 2023, Lemma A.8]
$W_{2}(\delta_{\bar{\theta}_{\star,\lambda}},\pi_{\lambda,\Theta}^{N})\leq\sqrt{\frac{d_{\theta}}{N\mu}},$ (17)
[…]
$W_{2}(\pi_{\lambda,\Theta}^{N},\mathcal{L}(\bm{\theta}_{n\gamma}^{N}))\leq e^{-\mu n\gamma}\bigg{(}\mathbb{E}[\|Z_{0}^{N}-z_{\star}\|^{2}]^{1/2}+\sqrt{\frac{d_{x}N+d_{\theta}}{N\mu}}\bigg{)}.$
Finally, the discretisation term is bounded by
@@
This bound is derived using a strategy similar to that used in classical Langevin methods.
@@ 4.4 Nonasymptotic analysis of PIPGLA
We introduce an extra assumption regarding the least norm element $\nabla^{0}g_{2}$ in the subdifferential of $g_{2}$, defined in A.4.
@@
For the convergence analysis of PIPGLA, we present a novel proof that differs from the error decomposition used for MYIPLA.
@@
Let A1, A2, A4 and C1 hold. Let $\theta_{n}^{N}$ denote the iterate (13) and $\bar{\theta}_{\star}$ be the maximiser of $p_{\theta}(y)$. Then for $\gamma\leq 1/L_{g_{1}}$ and $\gamma\leq\lambda\leq\gamma/(1-\mu\gamma)$, the following holds
$\mathbb{E}$ $\left[\|\theta_{n}^{N}-\bar{\theta}_{\star}\|^{2}\right]^{1/2}\leq\sqrt{\frac{\lambda^{n}(1-\gamma\mu)^{n}}{\gamma^{n}}}W_{2}(\mathcal{L}({Z_{0}^{N}}),\pi^{N})$
$+\sqrt{\frac{d_{\theta}}{N\mu}}+\sqrt{\frac{\lambda(2\gamma L_{g_{1}}(d_{\theta}+Nd_{x})+\lambda NC)}{N\left(1-\lambda(1-\mu\gamma)/\gamma\right)}},$
@@
See Appendix A.4 for the proof. Similarly to the previous result, we can split the errors as follows
@@
The concentration term $W_{2}(\delta_{\bar{\theta}_{\star}},\pi_{\Theta}^{N})$ can be bounded as in (17), thanks to the strong convexity assumption A1 and the fact that the set of non-differentiable points of the potential $U(\theta,x)$ has measure zero. For the other error term, we derive novel nonasymptotic bounds for general $\lambda$, unlike Salim et al. [2019] (see Corollary A.13 in App. A.4). Intuitively, $W_{2}(\pi_{\Theta}^{N},\mathcal{L}({\theta_{n}^{N}}))$ controls the convergence of the solution of the SDE to $\pi_{\Theta}^{N}$ and the error due to time discretisation. The bound for the latter term is derived without introducing the law of the solution of the SDE, $\mathcal{L}(\bm{\theta}_{n\gamma}^{N})$, in contrast to the approach taken for MYIPLA.